OpenStudy (anonymous):

find the quotient of 1 1/3 and 5/6.. Can you please show me the steps? thank u

OpenStudy (turingtest):

first convert 1 1/3 into a single fraction, what do you get?

OpenStudy (anonymous):

sorry 4/3

OpenStudy (turingtest):

good, so now we use the fact that a/b=a(1/b)\[{{4\over3}\over{5\over6}}={4\over3}*{6\over5}\]simplify and you have your answer.

OpenStudy (anonymous):

cancel first, multiply last

OpenStudy (anonymous):

\[\frac{4}{3}\times \frac{6}{5}=\frac{4}{1}\times \frac{2}{5}=\frac{8}{5}\]

OpenStudy (turingtest):

your answer is right too lamby, just simplify 9/15 by dividing top and bottom by three

OpenStudy (anonymous):

thank you both